OpenStudy (anonymous):

Please help! A cell phone company offers you a plan that charges a flat rate of $25 per month for 200 minutes of airtime. The rate of any time in excess of 200 minutes is $.12 per minute. A. Write an expression for the monthly dollar amount spent for the cell phone service if you want use x minutes of airtime in excess of 200 minutes. B. Write an inequality which indicates that you have budgeted no more than $40 per month to spend on the service. C. Determine the maximum number of minutes you can use in excess of 200 minutes so that you do not exceed your $40 per month budget.

jimthompson5910 (jim_thompson5910):

part A) we are told that "you want use x minutes of airtime in excess of 200 minutes", so automatically we know that x > 200 The company charges $25 per month for those 200 minutes. So that base charge of $25 is always there. Because x > 200, we're going to be charged $0.12 per minute on top of this $25 The remaining time over 200 minutes is x - 200. So 0.12(x-200) is added onto 25 to get 25 + 0.12(x-200) which is the total amount spent if x > 200 (ie you use more than 200 minutes)

jimthompson5910 (jim_thompson5910):

how would we do part B?

OpenStudy (anonymous):

Do we just double the equation?

OpenStudy (anonymous):

Wait no I don't think we do that.

jimthompson5910 (jim_thompson5910):

we need to write an inequality

jimthompson5910 (jim_thompson5910):

if T is the total cost, if we budget $40 for that total cost, then T <= 40 25 + 0.12(x-200) <= 40

jimthompson5910 (jim_thompson5910):

Notice how a) the total cost T is less than or equal to $40 and b) the total cost is also equal to 25 + 0.12(x-200)

jimthompson5910 (jim_thompson5910):

for part C, you need to solve the inequality found in part B

OpenStudy (anonymous):

So the answer would be 49 maximum number of minutes you can use in excess of 200 minutes so that you do not exceed your $40 per month budget?

jimthompson5910 (jim_thompson5910):

25 + 0.12(x-200) <= 40 25 + 0.12x-0.12*200 <= 40 25 + 0.12x-24 <= 40 0.12x+1 <= 40 0.12x <= 40-1 0.12x <= 39 x <= 39/0.12 x <= 325 So you can talk a max of 325 minutes to stay within budget of $40

jimthompson5910 (jim_thompson5910):

No you're not. You just need more practice.

OpenStudy (anonymous):

I haven't learned any of this and my teacher expect my classmates and I to do this at home.

jimthompson5910 (jim_thompson5910):

which part isn't making sense?

OpenStudy (anonymous):

I understand the equation part but when we have to solve the inequality is where I get lost.

jimthompson5910 (jim_thompson5910):

are you familiar with solving equations?

OpenStudy (anonymous):

Not really

jimthompson5910 (jim_thompson5910):

if you have something like x+5 = 10, then how do we isolate x?

OpenStudy (anonymous):

You subtract it from each side?

jimthompson5910 (jim_thompson5910):

yes, you undo the +5 and subtract it from both sides

jimthompson5910 (jim_thompson5910):

the same idea applies to 0.12x+1 <= 40 we undo the +1 by subtracting 1 from both sides

OpenStudy (anonymous):

Then whats the point of adding it there lol

jimthompson5910 (jim_thompson5910):

then we undo the 0.12 times x by dividing both sides by 0.12 to go from 0.12x <= 39 to x <= 39/0.12

jimthompson5910 (jim_thompson5910):

because the original equation allows you to find the total cost for any x

OpenStudy (anonymous):

Ohhhhh I understand.

OpenStudy (anonymous):

So then 39/0.12 gives us 325!

jimthompson5910 (jim_thompson5910):

and when you set that total cost equal to 40 and solve, you are essentially forcing the total cost to be $40 and then finding the x value that makes the total cost equal to 40

jimthompson5910 (jim_thompson5910):

exactly

OpenStudy (anonymous):

Oh thank you so much it's more clear to understand now.

jimthompson5910 (jim_thompson5910):

I'm glad it is